OpenStudy (anonymous):

precalc wiz?? you are in a hot air balloon and it has a 16 degree angle of elevation from a point 100 km away, then it rises to have an angle of elevation of 33 degrees from the same point, find how high it rose

OpenStudy (campbell_st):

I'll assume the rise is vertical  using sine rule \[\frac{h}{\sin17} = \frac{100}{\sin57}\] \[h = \frac{100\times \sin17}{\sin(57)}\] h = 34.86km strange answer,,,,